For f(x) = x2-5 and g(x) = sq rt of x, find h(x)=(fg)(x)
I get sq rt of x2-5. Is this correct?
: For f(x) = x2-5 and g(x) = sq rt of x, find h(x)=(fg)(x)
I get sq rt of x2-5. Is this correct?

Answer by midwood_trail(310)

Given f(x) = x^2 - 5 and g(x) = sqrt{x}, find h(x) = f(g(x)).
We plug the value of g(x) into f(x) and simplify, if possible.
By doing this, we will find h(x).
f(sqrt{x}) = (sqrt{x})^2 - 5
f(sqrt{x}) = x - 5
So, then h(x) = x - 5
Did you follow?